OpenStudy (anonymous):

how do i find the directrix and focus of this: (y-2)^2=8(x+1) please help! my test is tomorrow and I have no idea how to do it.

OpenStudy (anonymous):

do you know where the vertex of this parabola is ???

OpenStudy (anonymous):

do you if this is a parabola opening vertically or sideways??

OpenStudy (anonymous):

parabolas that open either up/down (vertically) have the equation: \(\large (y-k)^2=4p(x-h) \) parabolas that open either left/right (sideways) have the equation: \(\large (x-h)^2=4p(y-k) \) both parabolas have the vertex at (h, k). so once you know where the vertex is and which way the parabola opens, you can find the directrix to be p units away from the vertex.

OpenStudy (anonymous):

ok, so in this problem, the vertex is (-1,2)

OpenStudy (anonymous):

ok... and which way does this parabola open ????

OpenStudy (anonymous):

ummm....up? I know how to determine where it opens if I have the value of p. But I dont know how to find it in this problem.

OpenStudy (anonymous):

no.... if the squared term is the y, it is a parabola opening left/right... if the squared term is the x, it is a parabola opening up/down... which variable is squared in your equation??? the x or the y???

OpenStudy (anonymous):

y, so...it would be left of right...so, im guessing right?

OpenStudy (anonymous):

correct... now to determine if it opens to the left or to the right we have to determine p...

OpenStudy (anonymous):

ok. how?

OpenStudy (anonymous):

\(\large (y-2)^2=8(x+1) \) \(\large (y-2)^2=4\cdot \color {red}2(x+1) \) do you see that p=2 ???

OpenStudy (anonymous):

ok yeah

OpenStudy (anonymous):

and since p is positive, this parabola opens to the right....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, to find it I divide the coefficient by 4?

OpenStudy (anonymous):